# Self-Consistent Stochastic Finite-Temperature Modelling: Ultracold Bose Gases with Local (s-wave) and Long-Range (Dipolar) Interactions

### v1

@@ Abstract
We formulate a generalized self-consistent quantum kinetic theory including thermal fluctuations and stochastic contributions (arising from the leading order fluctuations around classical, deterministic dynamics) for modelling ultracold Bose gases interacting via a generic long-range interaction. Our generalised equations take the usual form of an effective field theory, separating coherent, low-lying, modes of the system from incoherent, higher-lying, thermal modes. The low-lying modes are described by a stochastic Langevin equation with two explicitly time-dependent collisional terms (corresponding to a dissipative and an energy-correcting contribution) and their corresponding additive and multiplicative stochastic noise terms. By coupling such an equation to an explicitly non-equilibrium gas of incoherent (thermal) particles described by a quantum Boltzmann equation, we thus extend beyond both earlier stochastic approaches (including the full stochastic projected Gross-Pitaevskii equation) and generalised kinetic models inspired by a two-gas picture (the so-called Zaremba-Nikuni-Griffin, or “ZNG” formalism) commonly used in the context of short-range interactions, such as those relevant in ultracold alkali atoms. Long-range interactions are further included into our model by the self-consistent addition of a Poisson-like equation for the long-range interaction potential, stimulated by our recent parallel work modelling dark matter as a combination of non-equilibrium (incoherent) particle [akin to cold dark matter] and (coherent) superfluid components [akin to fuzzy dark matter] [Proukakis et al., arXiv:2407.08860]. In the quantum gas context, our approach leads directly to a self-consistent model for finite-temperature Bose-Einstein condensation in a long-range interacting system within the regime where thermal fluctuations dominate over quantum fluctuations. While such an approach could be of general use for a variety of experimentally-accessible long-range interacting systems, we focus specifically here on the well-studied case of dipolar atomic condensates. In this particular context, we additionally supplement our Keldysh non-equilibrium analysis for fluctuations of the fast (incoherent) modes by a somewhat ad hoc extension of the slow (coherent) modes via the usual route of Bogoliubov-de Gennes equations: such treatment, which is known to be correct in the quantum-fluctuation-dominated regime, provides a mechanism for additionally including condensate fluctuations into our model. This enables us to postulate an extended dynamical model which may be suitable for interpolating between the regimes of quantum-dominated and thermal-dominated fluctuations, and to formulate an approximate criterion for the relative importance of such terms: nonetheless, we stress that – while our postulated extended model should provide an accurate quantitative description of the system in the two limiting cases – the extent of the applicability of such an extended model in their intermediate domain cannot be confirmed from first principles here, and remains a subject for future investigation.
@@ I Introduction
Modelling non-equilibrium quantum gases is a topic with a very rich history. While a significant part of the early literature was focused on obtaining equations for modelling the more strongly-interacting liquid Helium, the emergence of ultracold atomic gases in the 1990’s provided an ideal dilute, weakly-interacting, and strongly-controllable system for which ab initio theories could be directly relevant and compared to. Since then, established approaches [1, 2, 3, 4, 5] were revisited and appropriately extended, to take care of the novel features, observations and challenges presented by the dynamical richness introduced by the inhomogeneous external potentials. Mean-field equations, like the celebrated Gross-Pitaevskii Equation (GPE) [6, 7, 8, 9] first proposed as a suitable phenomenological model for liquid Helium, quickly became the workhorse of weakly-interacting ultracold quantum gases, modelling diverse scenarios [10, 11, 12], such as linear excitations, sound waves, bright and dark solitons, quantum vortices, Josephson dynamics, persistent currents and superflow, superfluid mixtures, spinor condensates, to mention a few. Although the GPE is in fact a remarkably versatile tool, both in terms of modelling near-zero-temperature weakly-interacting quantum gases – for which the quantum depletion is small – and also (perhaps somewhat counter-intuitively) the highly-populated modes of a finite-temperature system as an effective field theory [13, 14, 15, 16, 17], such a model – which is the simplest description based on a (non-relativistic) binary interaction Hamiltonian – was quickly supplemented by additional complexity as improved models better approximating the full system Hamiltonian gradually emerged over the past 3 decades (see, e.g., [18, 19, 20, 21, 22, 23, 24, 25, 26]).
@@
Loosely speaking, and beyond the addition of a phenomenological dissipation term first proposed by Pitaevskii [27, 28] which is being commonly applied as a first approximation for finite-temperature effects [29, 30], such finite-temperature approaches can be classified into a few rather different categories briefly summarized below. We note that we do not attempt to give here a fully comprehensive overview of all active approaches to date, for which we refer the reader to an earlier edited volume [18], related review articles [19, 20, 21] and works cited therein, or citing such works.
@@
The first broadly applicable method we wish to highlight here is based on techniques borrowed from the modelling of large quantum liquids, based on what might be called a ‘two-gas’ approximation, in direct analogy to the very successful ‘two-fluid’ model of liquid Helium [31, 32]. In the ‘two-gas’ model, one starts by directly invoking symmetry breaking of the Bose field operator, treating the beyond-condensate contribution perturbatively, under appropriate approximations. Such a methodology, first developed by Kirkpatrick and Dorfman [33, 34, 35, 36] (see also Eckern [37]) was significantly advanced by Zaremba, Nikuni and Griffin (‘ZNG’) [38, 23] and leads to a dissipative finite-temperature GPE self-consistently coupled to a quantum Boltzmann equation for the thermal cloud: such an approach has been successfully applied [39] to study damping of collective modes [40, 41, 42], solitons [43], vortices [44, 45, 46], Josephson junction dynamics [47, 48] and mixtures [49, 50, 51, 52, 53]. While an excellent model for a broad temperature range, and ideal for accurate treatment of collective modes, this model cannot describe the presence of large fluctuations, such as those arising in low-dimensional systems and near the critical region. As such, this model cannot describe the physics of phase transitions, although – rather remarkably – it can in fact capture fairly well quasi-adiabatic condensate growth once a well-formed small condensate (mimicked by an artifical initial seed) has formed [54, 55]. Another important shortcoming of the ZNG method manifesting itself in some regimes, and ultimately limiting its validity, is that it only accounts for single-particle ‘Hartree-Fock’ energies dressed by the potential and interactions, in the so-called ‘Popov’ approximation ignoring anomalous pair correlations [56], but it does not include quasiparticle physics associated with the Bogoliubov transformation [57] mixing single-particle creation and annihilation operators. This point will become relevant later on, when discussing the equations for modelling dipolar gases in the quantum droplet regime, when quantum corrections become critical [58].
@@
Various other closely-related approaches were developed in parallel to ZNG [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69], aiming to extend beyond such an approximation through the inclusion of so-called anomalous averages (see, e.g. , their review in Ref. [19]), but – despite limited applications [70] – these did not directly lead to broadly implemented numerical modelling (but see also discussion below). Nonetheless, we specifically highlight here the so-called (full) Hartree-Fock-Bogoliubov (HFB) equations, in which the normal and anomalous averages are defined by means of the coherence factors appearing within the Bogoliubov-de Gennes equations [19]: despite some problems associated with the decades-old discussion about conserving versus gapless approximations [71, 1, 63, 68, 69], the latter approach may have some benefits in our present context, as inclusion of Bogoliubov mixing actually encapsulates and can thus generate the condensate quantum depletion, the latter known to be critical to the dipolar droplet regime arising withing the modelling of dipolar condensates.
@@
Although the coupling to a quantum Boltzmann equation has already been discussed in [22, 105](see also [106, 107]), all numerical implementations of the various stochastic equations to date have made the explicit assumption of a large thermal cloud which remains essentially unperturbed during the evolution, and can thus act as a heat bath providing the dissipation and stochastic noise terms for the effective low-lying coherent region. In so doing, one is implicitly forced to drop the physically-exact relaxation requirement to a Bose-Einstein distribution, which instead becomes replaced by a Rayleigh-Jeans distribution, an excellent approximation for all highly-occupied modes (below the chosen cutoff), with typically many particles per mode, and at the characteristic temperatures studied [20]. Beyond Stoof’s pioneering reversible formation work [90], early dynamical studies on continuous atom laser operation [108] (revisited in  [109]), and rotating condensates [100], and a plethora of equilibrium studies [82, 85, 110, 86, 111, 112, 113, 87], this approach has been used to provide new insight into non-equilibrium phenomena in ultracold atomic gases, including most notably spontaneous defect generation during dynamical growth in the context of non-equilibrium phase transition physics and Kibble-Zurek universality [114, 115, 116, 117, 118, 119, 120, 121, 122, 123], and the subsequent relaxation process [124, 125, 126], soliton [127, 128, 129], vortex [130, 102, 104] and persistent current [118, 123, 131] dynamics. Other works include studies of collective mode dynamics [101, 42, 103], condensate mixtures [132, 133, 134, 135], and sound propagation [136]. Some subtleties associated with the above-mentioned approximations, including on the role of the ensemble, atom number conservation, and a different hybrid handling of the noise term which maintains a fully quantum description of the low-energy fields using the positive-P representation (which could prove crucial in the case of finite-temperature dipolar droplets) can be found in Refs. [137, 138, 139, 140]. See also Ref. [141, 142, 143] for other implementation in lower dimensions.
@@
As stated earlier, the primary aim of this introduction was simply to cast the discussion within this paper in the broader context of key preceeding quantum kinetic theory advances over the last few decades, such that our presented findings can be easily compared against such models. As such, we are aware of not having done full justice to other exciting methodological developments in the community, and we sincerely apologise for any unintentional omissions of alternative theories [^3] For completeness, and while not directly related to our present discussion despite their close connections, we also note here the existence of other approaches: most notably, the positive P-representation of quantum optics [152] for which the quantum operators $\hat{\Psi}$ and $\hat{\Psi}^{{\dagger}}$ are treated via stochastic equations for two independent classical fields $\phi$ and $\phi^{*}$ [153, 154], based on an underlying Fokker-Planck equation describing the positive-P quasiprobability phase space distribution function obtained from a master-equation approach; the field-theoretic two-particle-irreducible (2PI) formalism [155, 156, 157, 158]; the stochastic wavefunction approach [159]., or their implementations.
@@
Although the GPE has now become the workhorse model for weakly-interacting quantum gases, one should not forget that it was initially proposed for modelling liquid helium which is not well-described by simple contact interactions; in fact, a more accurate model of the helium interactions is based on a non-local interaction potential [160, 161, 21]. Although the familiar realm of dilute weakly-interacting ultracold atomic gases [10] is based on interactions which – under these conditions – can be effectively modelled as s-wave scattering, and thus a contact potential with an effective interaction strength set by the s-wave scattering length, the gallery of available quantum gases is constantly being broadened. In this context, we note the increasing interest in systems with non-negligible magnetic dipole moments, recently reviewed in Ref. [162]: these range from the well-studied dipolar condensates, which feature a combination of short-range (s-wave) and long-range (dipolar) interactions, to more elaborate experimentally-accessible systems, such as Rydberg atoms, or Bose-Einstein condensates of dipolar molecules, where particularly exciting progress has been made very recently [163]. With that in mind, the present work introduces in a very general manner a set of self-consistent finite-temperature equations describing partially-condensed systems exhibiting both short- and (arbitrary) long-range interactions, in a manner which includes both dissipative and stochastic noise terms, alongside self-consistently coupled Boltzmann equations for the incoherent particles and Poisson-like equations modelling the (arbitrary) long-range interaction. In the thermal-fluctuation-dominated regimes, such equations are more general than all previous treatments to date. Nonetheless – as we discuss below – the nature of our approximations is such that quantum fluctuations are not self-consistently treated to all orders, and the implications of – and potential remedies for – such an issue – are also touched upon later in the manuscript.
@@
In brief, this work is based on a standard Keldysh path integral formalism [164], which – using the tools discussed in detail in our companion paper [107] – leads to a self-consistent finite-temperature stochastic quantum model for ultracold Bose gases in the most general scenario of a co-existing short and long-range interaction between the particles, such as that found in dipolar gases. Our model generalizes beyond all above-discussed equations in the context of quantum gases, in a manner which will be clearly highlighted within the manuscript, where appropriate.
@@
Sec II briefly summarizes – following closely our related derivation in the cosmological setting [106, 107] – the key steps in the derivation of our main equations, which are presented for clarity in Sec. III. Secs. IV and V then explicitly relate such generalised equations to a broad range of applicable cold-atom theories discussed in the literature, in the specific contexts of local and dipolar interactions respectively. Our findings are briefly summarized in a final discussion section. Appendix A gives our full equations in a more general potential form, such that their mathematical expression can be readily available in the future for any more general non-local interaction potential becoming relevant. Appendix B gives some further context to the non-local operators appearing in the main manuscript, while Appendix C provides details on a parallel, but not self-consistent, inclusion of quantum effects to all orders in the near zero-temperature limit.
II System Hamiltonian and Effective Action for a Dipolar Atomic Gas
II.1 Basic Hamiltonian – Effective Action

### v2

# Dynamical Dipolar Condensate Finite Temperature Stochastic Gross–Pitaevskii–Boltzmann Model

@@ Abstract
We formulate a generalized self-consistent stochastic quantum kinetic theory for finite-temperature ultracold Bose gases interacting via a generic long-range interaction, applicable to a broad range of systems, by means of Keldysh non-equilibrium field theory: such model is explicitly cast in the context of dipolar atomic gases, and is also shown to encompass established stochastic and kinetic treatments for ultracold atomic gases with local interactions as special cases. The condensate and low-lying modes are collectively described by a stochastic Gross-Pitaevskii equation with two collisional terms and their corresponding stochastic noise terms, with thermal particles dynamically modelled through a self-consistently coupled Quantum Boltzmann equation and dipolar interactions included by means of a coupled Poisson-like equation. Additional use of Bogoliubov-de Gennes analysis generating the Lee-Huang-Yang correction term relevant in the $T=0$ quantum-fluctuation-dominated regime, allows us to postulate the extension of such model offering a plausible scheme for interpolating between quantum-dominated and thermal-dominated fluctuation regimes, the consistency of which remains to be tested against experimental observations.
@@ I Introduction
The aim of this work is to discuss a generalized framework unifying some of the most common approaches for modelling ultracold atomic gases with contact interactions, and extend it to model systems with arbitrary long-range interactions at finite temperatures. While such a model can be generically relevant for a broad range of physical systems, the emphasis in this work is on the modelling of dipolar atomic gases [1, 2], due to the significant experimental impetus in this field (see, e.g., [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]), and the current lack of a fully self-consistent dynamical theory encompassing both quantum and thermal effects. We believe that this work [^1] Note that the full model presented in this work was first posted on arXiv in July 2024, accessible as arXiv:2407.20178v1. makes significant steps in that direction, offering an extended complementary view to existing approaches for dipolar gases, which primarily focus on the role of quantum fluctuations near the instability regime around $T\approx 0$ where quantum fluctuations (and the so-called Lee-Huang-Yang [21] correction to the ground state energy) become important [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 12, 32, 8, 33, 16, 34, 35], or $T>0$ settings either in equilibrium [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], by means of quantum Monte Carlo [51, 52, 53], or based on alternative (often simpler) dynamical models [22, 54, 55, 56, 57, 58, 59]. Implications of this model and the quantification of the relevance of the various terms to different experimentally relevant scenarios (including different extent of near- and far-from-equilibrium, and different dimensionalities) remain outside the scope of the present work and are deferred to future work.
This paper is structured as follows: Sec. II gives a brief historical overview of some of the most common methods used in the field of ultracold atomic gases interacting via the usual s-wave scattering. Sec. III gives the basic system Hamiltonian, and outlines the key steps in the derivation of our model in the thermal-dominated regime, with our final equations for dipolar condensates1 given in Sec. IV, with corresponding equations in a generic long-range potential given in Appendix A. Sec. V shows that in the limit of only s-wave scattering (contact interactions) our equations constitute an extended hybridization of the commonly used Zaremba-Nikuni-Griffin (‘ZNG’) and Stochastic (Projected) Gross-Pitaevskii Equation (S(P)GPE) established formalisms, clearly highlighting the steps for reducing our theory to such simpler, yet remarkably powerful, models. In Sec. VI we focus on dipolar atomic gases, comment on the interplay between thermal and quantum fluctuations, and – by analogy to established models for contact-interacting cold atoms – we provide reduced sets of equations for dipolar equations valid in appropriate limits. We anticipate that equations of this kind will be extensively used in the following decade to address a number of problems of direct relevance to ultracold dipolar gases. Appendices B and C provide further technical details associated with the operator appearing in our Poisson-like equation for the dipolar interaction and the emergence of the Lee-Huang-Yang contribution to the system’s ground state energy.
II Brief Historical Overview of Dynamical Bose Gas Theories beyond the GPE
@@
Modelling non-equilibrium quantum gases is a topic with a very rich history. While a significant part of the early literature was focused on obtaining equations for modelling the more strongly-interacting liquid Helium, the emergence of ultracold atomic gases in the 1990’s provided an ideal dilute, weakly-interacting, and strongly-controllable system for which ab initio theories could be directly relevant and compared to. Since then, established approaches [60, 61, 62, 63, 64] were revisited and appropriately extended, to take care of the novel features, observations and challenges presented by the dynamical richness introduced by the inhomogeneous external potentials. Mean-field equations, like the celebrated Gross-Pitaevskii Equation (GPE) first proposed as a suitable phenomenological model for liquid Helium [65, 66, 67, 68], quickly became the workhorse of weakly-interacting ultracold quantum gases, modelling diverse scenarios [69, 70, 71], such as linear excitations, sound waves, bright and dark solitons, quantum vortices, Josephson dynamics, persistent currents and superflow, superfluid mixtures, spinor condensates, to mention a few. The GPE is in fact a remarkably versatile tool, both in terms of modelling near-zero-temperature weakly-interacting quantum gases – for which the quantum depletion is small – and also (perhaps somewhat counter-intuitively) the highly-populated modes of a finite-temperature system as an effective field theory [72, 73, 74, 75, 76, 77]. Such a model – which is the simplest classical description based on a (non-relativistic) binary interaction Hamiltonian – gradually evolved into more extended models, such as the ‘Zaremba-Nikuni-Griffin’ (ZNG) and the Stochastic (Projected) Gross-Pitaevskii Equation (SPGPE), as alternative models of the full system Hamiltonian – see, for example the recent reviews  [78, 79, 80, 81, 82, 83, 84, 85].
@@
Beyond the addition of a phenomenological dissipation term first proposed by Pitaevskii [86, 87] which is being commonly applied as a first approximation for finite-temperature effects (see, e.g., [88, 89]), such finite-temperature approaches can be loosely speaking classified into a few rather different categories briefly summarized below. We feel it is useful to do so here, in order to motivate the different limits of dipolar gas theories proposed in Sec. VI, but note that we do not attempt to give a fully comprehensive overview of all active approaches to date, for which we refer the reader to an earlier edited volume [78], related review articles [79, 80, 81] and works cited therein, or citing such works.
@@
The first broadly applicable method we wish to highlight here is based on techniques borrowed from the modelling of macroscopic quantum liquids, based on what might be called a ‘two-gas’ approximation, in direct analogy to the very successful ‘two-fluid’ model of liquid Helium [90, 91]. In the ‘two-gas’ model, one starts by directly invoking symmetry breaking of the Bose field operator, treating the beyond-condensate contribution perturbatively, under appropriate approximations. Such a methodology, first developed by Kirkpatrick and Dorfman [92, 93, 94, 95] (see also Eckern [96]) was significantly advanced by Zaremba, Nikuni and Griffin (‘ZNG’) [97, 83] and leads to a dissipative finite-temperature GPE self-consistently coupled to a quantum Boltzmann equation for the thermal cloud: such an approach has been successfully applied [98] to study damping of collective modes [99, 100, 101], solitons [102], vortices [103, 104, 105], Josephson junction dynamics [106, 107] and mixtures [108, 109, 110, 111, 112]. While an excellent model for a broad temperature range, and ideal for accurate treatment of collective modes, this model cannot describe the presence of large fluctuations, such as those arising in low-dimensional systems and near the critical region. As such, this model cannot describe the physics of phase transitions, although – rather remarkably – it can in fact capture fairly well quasi-adiabatic condensate growth once a well-formed small condensate (mimicked by an artifical initial seed) has formed [113, 114]. Another important shortcoming of the ZNG method manifesting itself in some regimes, and ultimately limiting its validity, is that it only accounts for single-particle ‘Hartree-Fock’ energies dressed by the potential and interactions, in the so-called ‘Popov’ approximation ignoring anomalous pair correlations [115], but it does not include quasiparticle physics associated with the Bogoliubov transformation [116] mixing single-particle creation and annihilation operators. This point will become relevant later on, when discussing the equations for modelling dipolar gases in the quantum droplet regime, when such corrections become critical [117].
@@
Various other closely-related approaches were developed in parallel to ZNG [118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128], aiming to extend beyond such an approximation through the inclusion of so-called anomalous averages (see, e.g., their review in Ref. [79]), but – despite limited applications [129] – these did not directly lead to broadly implemented numerical modelling (but see also discussion below). Nonetheless, we specifically highlight here the so-called (full) Hartree-Fock-Bogoliubov (HFB) equations, in which the normal and anomalous averages are defined by means of the coherence factors appearing within the Bogoliubov-de Gennes equations [79]: despite some problems associated with the decades-old discussion about conserving versus gapless approximations [130, 60, 122, 127, 128], the latter approach may have some benefits in our present context, as inclusion of Bogoliubov mixing actually encapsulates and can thus generate the condensate quantum depletion, the latter known to be critical to the dipolar droplet and supersolid regimes.
@@
Although the coupling to a quantum Boltzmann equation has already been discussed in [82, 164](see also [165, 166]), all numerical implementations of the various stochastic equations to date have made the explicit assumption of a large thermal cloud which remains essentially unperturbed during the evolution, and can thus act as a heat bath providing the dissipation and stochastic noise terms for the effective low-lying coherent region. In so doing, one is implicitly forced to drop the physically-exact relaxation requirement to a Bose-Einstein distribution, which instead becomes replaced by a Rayleigh-Jeans distribution [but see also Ref. [167]], an excellent approximation for all highly-occupied modes (below the chosen cutoff), with typically many particles per mode, and at the characteristic temperatures studied [80]. Beyond Stoof’s pioneering reversible formation work [149], early dynamical studies on continuous atom laser operation [168] (revisited in  [169]), and rotating condensates [159], and a plethora of equilibrium studies [141, 144, 170, 145, 171, 172, 173, 146], this approach has been used to provide new insight into non-equilibrium phenomena in ultracold atomic gases, including most notably spontaneous defect generation during dynamical growth in the context of non-equilibrium phase transition physics and Kibble-Zurek universality [174, 175, 176, 177, 178, 179, 180, 181, 182, 183], and the subsequent relaxation process [184, 185, 186], soliton [187, 188, 189], vortex [190, 161, 163] and persistent current [178, 183, 191] dynamics. Other works include studies of collective mode dynamics [160, 101, 162], condensate mixtures [192, 193, 194, 195], and sound propagation [196]. Some subtleties associated with the above-mentioned approximations, including on the role of the ensemble [167], atom number conservation, and a different hybrid handling of the noise term which maintains a fully quantum description of the low-energy fields using the positive-P representation (which could prove crucial in the case of finite-temperature dipolar droplets) can be found in Refs. [197, 198, 167, 199]. See also Ref. [200, 201] for other implementation in lower dimensions (and the somewhat related work [202].
@@
As stated earlier, the primary aim of this introduction was simply to cast the discussion within this paper in the broader context of key preceeding quantum kinetic theory advances over the last few decades, such that our presented findings can be easily compared against such models. As such, we are aware of not having done full justice to other exciting methodological developments in the community, and we sincerely apologise for any unintentional omissions of alternative theories [^4] For completeness, and while not directly related to our present discussion despite their close connections, we also note here the existence of other approaches: most notably, the ‘raw’ truncated Wigner for the whole field [206, 209, 211], the positive P-representation of quantum optics [212] for which the quantum operators $\hat{\Psi}$ and $\hat{\Psi}^{{\dagger}}$ are treated via stochastic equations for two independent classical fields $\phi$ and $\phi^{*}$ [206, 213], based on an underlying Fokker-Planck equation describing the positive-P quasiprobability phase space distribution function obtained from a master-equation approach; the field-theoretic two-particle-irreducible (2PI) formalism [214, 215, 216, 217]; the stochastic wavefunction approach [218]., or their implementations.
@@
Although the GPE has now become the workhorse model for weakly-interacting quantum gases, one should not forget that it was initially proposed for modelling liquid helium which is not well-described by simple contact interactions; in fact, a more accurate model of the helium interactions is based on a non-local interaction potential [219, 220, 81]. Although the familiar realm of dilute weakly-interacting ultracold atomic gases [69] is based on interactions which – under these conditions – can be effectively modelled as s-wave scattering, and thus a contact potential with an effective interaction strength set by the s-wave scattering length, the gallery of available quantum gases is constantly being broadened. In this context, we note the increasing interest in systems with non-negligible magnetic dipole moments, recently reviewed in Ref. [1]: these range from the well-studied dipolar condensates, which feature a combination of short-range (s-wave) and long-range (dipolar) interactions, to more elaborate experimentally-accessible systems, such as Rydberg atoms, or Bose-Einstein condensates of dipolar molecules, where particularly exciting progress has been made very recently [221]. With that in mind, the present work introduces in a very general manner a set of self-consistent finite-temperature equations describing partially-condensed systems exhibiting both short- and (arbitrary) long-range interactions, in a manner which includes both dissipative and stochastic noise terms, alongside self-consistently coupled Boltzmann equations for the incoherent particles and Poisson-like equations modelling the (arbitrary) long-range interaction. In the thermal-fluctuation-dominated regimes, such equations are more general than all previous treatments to date. Nonetheless – as we discuss below – the nature of our approximations is such that quantum fluctuations are not self-consistently treated to all orders, and the implications of – and potential remedies for – such an issue – are also touched upon within this manuscript.
@@
In brief, this work is based on a standard Keldysh path integral formalism [222], which – using the tools discussed in detail in our companion paper [166] – leads to a self-consistent finite-temperature stochastic quantum model for ultracold Bose gases in the most general scenario of a co-existing short and long-range interaction between the particles, such as that found in dipolar gases. Our model thus generalizes beyond the above-discussed quantum gas models, as highlighted (where appropriate) within the manuscript.
@@
Sec III briefly summarizes – following closely our related derivation in the context of gravitational coupling [165, 166] – the key steps in the derivation of our main equations, which are presented for clarity in Sec. IV. Secs. V and VI then explicitly relate such generalized equations to a broad range of applicable cold-atom theories discussed in the literature, in the specific contexts of local and dipolar interactions respectively. Our findings are briefly summarized in a final discussion section. Appendix A gives our full equations in a more general potential form, such that their mathematical expression can be readily available in the future for any more general non-local interaction potential becoming relevant. Appendix B gives some further context to the non-local operators appearing in the main manuscript, while Appendix C provides details on the inclusion of quantum effects to all orders in the near zero-temperature limit.
III System Hamiltonian and Effective Action for a Dipolar Atomic Gas
III.1 Basic Hamiltonian – Effective Action